# Rationale-Enhanced Decoding for Multi-modal Chain-of-Thought

### v1

@@ 4.4 Intervention Analysis
We examine whether RED can overcome the issue of multi-modal CoT not being able to leverage rationale, as described in Section 2.3. We hypothesized that the performance should improve with higher-quality rationales and degrade with irrelevant ones. To evaluate this, we conduct intervention analysis using GPT-4 Achiam_2023_GPT4 generated rationales as the high-quality rationale, and the randomly swapped irrelevant rationales in the same way as Section 2.3. Table 4 shows the results of this intervention analysis on GQA with CCoT, where Self denotes using the rationales generated from the backbone LVLM itself, GPT-4 denotes using high-quality rationales generated by GPT-4, and Random denotes using irrelevant rationales swapped from other instances. CCoT + RED with GPT-4 rationales greatly improved performance, Random rationales degraded it. The substantial performance gains observed with CCoT + RED using GPT-4 rationales, exceeding the CCoT counterparts, indicate that RED effectively leverages improved rationale content for multi-modal reasoning. These results indicate not only that RED can overcome the issue of conventional multi-modal CoT reasoning but also that RED offers some interpretability in the reasoning process, which is dependent on the rationale.
@@
Table 4: Intervention analysis on GQA with CCoT. Performance is shown when using self-generated rationales (Self), rationales generated by GPT-4 (GPT-4), and randomly swapped irrelevant rationales (Random).
Figure 2: Performance trends on GQA accuracy with increasing LVLM parameter sizes using Baseline, CCoT, CCoT + RED for Gemma-3 and Qwen2.5-VL families.
@@ 4.5 Scalability for Larger LVLMs
We test RED on LVLMs with a larger parameter size for evaluating the scalability. We additionally utilized Gemma-3-27B, Qwen2.5-VL-32B/-72B. These are increased language model sizes, and the vision encoders remain unchanged. Figure 2 shows the scalability on larger LVLMs. Contrary to intuition, neither Baseline nor CCoT consistently improved performance with model size. This is potentially because, as shown in Section 2.3, the decoding process often over-relies on visual input tokens at the expense of rationales; a similar observation is found in Al-tahan_NeurIPS24_unibench . In contrast, our RED consistently improved performance in proportion to model size, indicating that RED can leverage sophisticated rationales from larger language models. This implies that RED could be a key factor in unlocking the full potential of larger LVLMs by enabling more effective multi-modal CoT.
@@
4.6 Qualitative Evaluation
Figure 3 shows qualitative examples from GQA; we performed reasoning with CoT and CCoT at the top and bottom of the figure, respectively. While generated rationales points to the correct answer, naive CoT/CCoT produce incorrect responses, highlighting their failure to leverage the rationales by $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$. Our RED, in contrast, successfully extracts the conclusion and related attributes from the rationales in both cases.
5 Related Work
Large Vision Language Models (LVLMs).
Large vision language models (LVLMs) integrate visual encoders with LLMs, often via alignment training to represent images in the LLM input space liu_NeurIPS23_llava ; Dai_NeurIPS23_InstructBLIP ; Zhu_ICLR24_MiniGPT ; Chen_NeurIPS24_InternVL . Despite remarkable multi-modal capabilities Laurenccon_NeurIPS24_matters_of_lvlms ; Parekh_NeurIPS24_conceptual_explainability_for_lvlms ; Neo_ICLR25_interpetting_lvlms , LVLMs face challenges like poor visual recognition Zhang_NeurIPS24_why_lvlm_bad_classification , object hallucination Li_EMNLP23_evaluating_lvlm ; Gunjal_AAAI24_detecting_object_hallucination_lvlm ; Wu_ICML24_analyzing_hallucination_lvlm , and misalignment between image and text tokens Zeng_CVPR24_investigating_compositionality_challenge_lvlm ; Wang_NeurIPS24_picture_worth_lvlm ; Campbell_NeurIPS24_understanding_binding_limits_lvlms . Furthermore, as our preliminary experiments (Section 2.3) highlight, LVLMs also struggle to effectively leverage rationales in CoT reasoning. Previous solutions include preference/reward tuning Zhou_NeurIPS24_calibrated_reward_tuning ; Sun_ACL24F_aligning_lvlm_augmented_rlhf , improved visual instruction tuning Ossowski_ACL24_olive ; Chen_ICLR25_perturbollava ; Shi_ICLR25_enhancing_cognition_lvlm ; Chow_ICLR24_unified_training_lvlm ; Sirnam_ECCV24_x-former , auxiliary model-enhanced decoding Yu_ECCV24_attention_prompting ; Zhang_ICLR25_selfcorrecting_generative_decoding_lvlm ; Sun_ICLR25_fast_slow_visual_agents , plug-and-play decoding Manevich_ACL24_LCD ; Leng_CVPR24_VCD ; Wang_ACL24F_ICD ; Ghosh_ICLR25_VDGD , and CoT prompting Zhang_TMLR_mm-cot ; Zheng_NeurIPS23_ddcot ; Mondal_AAAI24_kam-cot ; He_AAAI24_mm_latent_cot ; Zheng_LREC24_enhancing_mmcot_soft_negative_sampling ; Mitra_CVPR24_CCoT . Our RED is categorized into a plug-and-play decoding for improving CoT prompting. While many CoT prompting methods focus on improving rationale generation via specific prompting or auxiliary model training (see Section 2.2), they often still rely on standard decoding with $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ that may not leverage rationales. Our work, RED, diverges by introducing a novel plug-and-play decoding strategy designed to enhance rationale utilization in CoT reasoning, without additional training or auxiliary models.
Plug-and-play Decoding Strategies for LVLMs.
Plug-and-play decoding strategies Manevich_ACL24_LCD ; Leng_CVPR24_VCD ; Wang_ACL24F_ICD ; Ghosh_ICLR25_VDGD are relevant as they operate at inference time without additional training like RED. These decoding strategies are mainly focused on mitigating object hallucination by contrastive decoding Li_ACL23_contrastive_decoding . For instance, LCD Manevich_ACL24_LCD contrasts $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ with $p_{\theta}(y_{i}|\bm{y}_{<i},q)$ to mitigate the language prior effects; VCD Leng_CVPR24_VCD subtracts hallucinated predictions by contrasting $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ with $p_{\theta}(y_{i}|\bm{y}_{<i},x^{\prime},q)$ ($x^{\prime}$ is the corrupted image). Thus, these decoding strategies basically aim to modulate image-conditional probability $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ to mitigate object hallucination. However, these methods do not ensure LVLMs faithfully use CoT rationales; they refine image-conditional output, not harmonize it for CoT reasoning. Therefore, RED is complementary to these decoding strategies because it grounds predictions by multiplying distinct image-conditional $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$, which is potentially pre-modulated by other methods, and rationale-conditional $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$. This allows combining RED with other plug-and-play methods for synergistic benefits of better hallucination mitigation and rationale grounding.
@@ Appendix B Additional Experiments
As alternatives for Eq. (8), we used (i) rationale-conditional probability $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$, (ii) mixture-of-experts $(1-\lambda)p_{\theta}(y_{i}|\bm{y}_{<i},x,q)+\lambda p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$, and (iii) reversed power-of-experts $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)\times p_{\theta}(y_{i}|\bm{y}_{<i},x,q)^{\lambda}$. We see that our formulation by Eq. 8 achieves the best accuracy. While rationale-conditional probability $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$ is grounded on the rationale $r$ in the next token prediction, it may cause the over-reliance on textual information, as reported in [42]. For mixture-of-experts, it produces the next tokens by the sum of $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ and $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$, which may fail to effectively collaborate image- and rationale-conditional prediction because it prioritizes tokens that have a high probability assigned to one of the distributions, similar to an OR operation. The result of reverse power-of-experts demonstrates the importance of maximizing the rationale-conditional probability as a reward in Eq. (7). Since the effects of the image $x$ in $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ is already dominant as discussed in Section 2.3, predicting by $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)\times p_{\theta}(y_{i}|\bm{y}_{<i},x,q)^{\lambda}$ further weaken the effects of the rationale $r$, leading a suboptimal performance similar to Baseline and CCoT. In contrast, our formulation of RED can appropriately focus on tokens important for both the image- and rationale-conditional probabilities by the multiplication, like an AND operation. More importantly, Eq. (8) has the theoretical guarantee that maximizes the rationale-conditional probability while maintaining the image-conditional prediction capability, achieving more stable and better accuracy gain for multi-modal CoT prompting.
@@ B.3 Effects of $\lambda$
Appendix C Reproducibility Statement
As described in Section 4.1, the implementation of the proposed method uses a publicly available code base. The LVLMs backbones are publicly available on the HuggingFace repositories (e.g., Gemma-3: https://huggingface.co/google). All datasets are also available on the web; see the references in Section 4.1 for details. For the computation resources, we used a 24-core Intel Xeon CPU with an NVIDIA H100 GPU with 80GB VRAM. We will make the code available on the public repository if the paper is accepted.
Appendix D Broader Impacts

### v2

@@ 5.4 Intervention Analysis
We examine whether RED can overcome the issue of multi-modal CoT not being able to leverage rationale, as described in Section 3.3. We hypothesized that the performance should improve with higher-quality rationales and degrade with irrelevant ones. To evaluate this, we conduct intervention analysis using GPT-4 [1] generated rationales as the high-quality rationale, and the randomly swapped irrelevant rationales in the same way as Section 3.3. Table 4 shows the results of this intervention analysis on GQA with CCoT, where Self denotes using the rationales generated from the backbone LVLM itself, GPT-4 denotes using high-quality rationales generated by GPT-4, and Random denotes using irrelevant rationales swapped from other instances. Note that we used the same $\lambda$ tuned with the Self rationales across all cases. CCoT + RED with GPT-4 rationales greatly improved performance, Random rationales degraded it. The substantial performance gains observed with CCoT + RED using GPT-4 rationales, exceeding the CCoT counterparts, indicate that RED effectively leverages improved rationale content for multi-modal reasoning. These results indicate not only that RED can overcome the issue of conventional multi-modal CoT reasoning but also that RED offers some interpretability in the reasoning process, which is dependent on the rationale.
@@
Table 4: Intervention analysis on GQA with CCoT. Performance is shown when using self-generated rationales (Self), rationales generated by GPT-4 (GPT-4), and randomly swapped irrelevant rationales (Random). RED can be improved by a higher-quality rationale and degraded by a lower-quality one, demonstrating its faithfulness in grounding on rationales.
Figure 3: Performance trends on GQA accuracy with increasing LVLM parameter sizes (Baseline, CCoT, CCoT + RED for Gemma-3 and Qwen2.5-VL families). RED can consistently improve Baseline and CCoT in any models, and can unlock further performance scalability according to model sizes.
@@ 5.5 Scalability for Larger LVLMs
We test RED on LVLMs with a larger parameter size for evaluating the scalability. We additionally utilized Gemma-3-27B, Qwen2.5-VL-32B/-72B. These are increased language model sizes, and the vision encoders remain unchanged. Figure 3 shows the scalability on larger LVLMs. Contrary to intuition, neither Baseline nor CCoT consistently improved performance with model size. This is potentially because, as shown in Section 3.3, the decoding process often over-relies on visual input tokens at the expense of rationales; a similar observation is found in [2]. In contrast, our RED consistently improved performance in proportion to model size, indicating that RED can leverage sophisticated rationales from larger language models. This implies that RED could be a key factor in unlocking the full potential of larger LVLMs by enabling more effective multi-modal CoT.
5.6 Qualitative Evaluation
Figure 4 shows qualitative examples from GQA; we performed reasoning with CoT and CCoT at the top and bottom of the figure, respectively. While generated rationales point to the correct answer, naive CoT/CCoT produce incorrect responses, highlighting their failure to leverage the rationales by $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$. In contrast, RED extracts the conclusion and related attributes from the rationales in both cases.
5.7 Inference Efficiency
We evaluate the inference efficiency of RED. The inference steps are decomposed into (i) rationale generation by Eq. (5) and (ii) answer generation by Eq. (8). The former is the same as the existing CoT, and the latter requires a dual-forward pass, which we implemented using batch-parallel decoding. To assess the latency, we measured the average inference time for each query across all benchmarks with Gemma-3-12B on Baseline, CoT, and RED. The results were 3.01, 5.27, and 5.05 for Baseline, CoT, and RED, respectively. RED was faster than CoT because the contexts of $(x,q)$ and $(r,q)$ of the batch-parallel decoding were shorter than the full context $(x,r,q)$ of CoT, emphasizing the efficiency of our method. Nonetheless, slightly increasing GPU memory consumption due to parallel RED is a limitation that should be resolved in future work.
@@ Appendix B Additional Experiments
As alternatives for Eq. (8), we used (i) rationale-conditional probability $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$, (ii) mixture-of-experts $(1-\lambda)p_{\theta}(y_{i}|\bm{y}_{<i},x,q)+\lambda p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$, and (iii) reversed power-of-experts $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)\times p_{\theta}(y_{i}|\bm{y}_{<i},x,q)^{\lambda}$. We see that our formulation by Eq. (8) achieves the best accuracy. While the rationale-conditional probability $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$ is grounded on the rationale $r$ in the next token prediction, it may cause the over-reliance on textual information, as reported in [58]. For mixture-of-experts, it produces the next tokens by the sum of $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ and $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$, which may fail to effectively collaborate image- and rationale-conditional prediction because it prioritizes tokens that have a high probability assigned to one of the distributions, similar to an OR operation. The result of reverse power-of-experts demonstrates the importance of maximizing the rationale-conditional probability as a reward in Eq. (7). Since the effects of the image $x$ in $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ is already dominant as discussed in Section 2.3, predicting by $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)\times p_{\theta}(y_{i}|\bm{y}_{<i},x,q)^{\lambda}$ further weaken the effects of the rationale $r$, leading a suboptimal performance similar to Baseline and CCoT. In contrast, our formulation of RED can appropriately focus on tokens important for both the image- and rationale-conditional probabilities by the multiplication, like an AND operation. More importantly, Eq. (8) has the theoretical guarantee that maximizes the rationale-conditional probability while maintaining the image-conditional prediction capability, achieving more stable and better accuracy gain for multi-modal CoT prompting.
@@ B.3 Effects of $\lambda$
B.4 Evaluations on MMMU/MMMU-Pro
Table 7 shows the results on the MMMU benchmarks. RED largely outperformed baselines, indicating that it remains effective on more complex multi-modal reasoning tasks.
Table 7: Evaluations on MMMU/MMMU-Pro with Qwen2.5-VL-7B
MMMU MMMU-Pro
Model Accuracy (%) Speed (ms/token) Accuracy (%) Speed (ms/token)
Baseline 50.3 9.5 35.2 12.4
VCD + ICD 48.5 12.1 33.1 14.7
CoT 55.5 17.5 37.2 28.7
CoT + RED (Ours) 61.6 14.5 40.5 16.2
Appendix C Broader Impacts